OpenStudy (anonymous):

you are expecting a call anytime between 3:00pm and 5:pm. at 3:20 you discover that someone laft the phone off the hook. what is the probability that you missed the call?

OpenStudy (anonymous):

\[\frac{20}{120}\]

OpenStudy (anonymous):

better known as \[\frac{1}{6}\] because the phone was off the hook for \[\frac{1}{6}\text{th}\] of the time

OpenStudy (anonymous):

about what would that be in %?